# Optimal Rates for Generalization of Gradient Descent for Deep ReLU Classification

### v1

@@ 4.2 Generalization Analysis
In previous work (Chen et al., 2021), they derived the bound of $\widetilde{O}\big(4^{L}L\sqrt{{mF(\overline{\mathbf{W}})}/{n}}\big)$. Specifically, they use the generalization analysis in Bartlett et al. (2017), which requires to estimate the following term
@@
To bound the term $G^{\prime}=\sup_{z,\mathbf{W}\in\mathcal{W}_{1}}\ell(yf_{\mathbf{W}}(\mathbf{x}))$, we analyze the difference $f_{\mathbf{W}}(\mathbf{x})-f_{\overline{\mathbf{W}}}(\mathbf{x})$. Since both $\mathbf{W},\overline{\mathbf{W}}\in\mathcal{B}_{R}(\mathbf{W}(0))$ for some $R$, we only need to study the local variation $f_{\mathbf{W}}(\mathbf{x})-f_{\mathbf{W}(0)}(\mathbf{x})$. This approach necessitates characterizing the uniform behavior of deep networks in $\mathcal{B}_{R}(\mathbf{W}(0))$, specifically establishing control over their Lipschitz constants near initialization. Existing works usually lead to an exponential dependence on $L$ (Xu and Zhu, 2024; Taheri et al., 2025), thus resulting in a $e^{O(L)}$ term in the generalization bound. In particular, Lemma F.3 and F.5 in Liu et al. (2020) pointed out that $\|h^{l}(\mathbf{x})\|\leq C^{L},\|\partial f_{\mathbf{W}}(\mathbf{x})/\partial h^{l}(\mathbf{x})\|_{2}\leq C^{L-l+1}\sqrt{m}$. Based on these observations, Taheri et al. (2025) showed that
@@
reducing our task to bounding the hidden layer perturbation. Previous approaches, including Xu and Zhu (2024) and Du et al. (2019b), employ a recursive estimation:
@@
In contrast to previous work, we develop the covering-number strategy to avoid the exponential dependence on depth. Specifically, we first show that for $n$ training data, if $m=\widetilde{\Omega}(L^{10}\log(n)R^{2})$, then $\|h^{l}(\mathbf{x}_{i})-h^{l}_{0}(\mathbf{x}_{i})\|_{2}=\widetilde{O}\left(\frac{L^{2}R}{\sqrt{m}}\right)$ holds for $i\in[n],l\in[L]$ (Lemma 15). We further take a $1/(C^{L}\sqrt{m})$-covering $D=\{\mathbf{x}^{j}:j=1,\ldots,|D|\}$ of the input space. Recall that the input space $\mathcal{X}=S^{d-1}$, it is well known from Corollary 4.2.13 in Vershynin (2018) that the number of $1/(C^{L}\sqrt{m})$-covering is given by $|D|\leq(1+2C^{L}\sqrt{m})^{d}$. Applying Lemma 15 to $D$ derives that if $m=\widetilde{\Omega}(L^{10}\log(|D|)R^{2})$
@@ 4.3 Optimal rates on NTK-separable data
In this section, we apply our general analysis to NTK-separable data (Ji and Telgarsky, 2020; Nitanda et al., 2020; Chen et al., 2021; Taheri and Thrampoulidis, 2024; Deora et al., 2023), and obtain the optimal rates.
@@
This means that the dataset is separable by the NTK feature at initialization with a margin $\gamma$. Nitanda et al. (2020) pointed out that this assumption is weaker than positive eigenvalues of NTK Gram matrix, which has been widely used in the literature (Du et al., 2019b, a; Arora et al., 2019a). With the above assumption, we have the following optimal risk bound on NTK separable data. The proof is given in Appendix D.
@@
Under Assumption 3, Theorem 3.3 and Proposition 4.2 of Chen et al. (2021) show that when the network width satisfies $m=\widetilde{\Omega}(L^{22}/\gamma^{8})$, the training error is of the order $\widetilde{O}(1/T)$. We achieve the same guarantee under a significantly milder width condition of $m=\widetilde{\Omega}(L^{16}/\gamma^{8})$. This improvement is enabled by two key technical advances: a sharper bound for (4), and a tighter estimate of the iterate distance $\|\mathbf{W}(t)-\overline{\mathbf{W}}\|_{F}$. Specifically, we improve the bound in (4) by a factor of $L^{1/3}$, and show that $\|\mathbf{W}(t)-\overline{\mathbf{W}}\|_{F}=\widetilde{O}(1/\gamma)$, improving upon the previous $\widetilde{O}(\sqrt{L}/\gamma)$ bound. Together, these refinements reduce the required network width by a factor of $L^{6}$.
@@
Under similar assumptions, Ji and Telgarsky (2020) derived the bound $\widetilde{O}\left(\frac{1}{\gamma^{2}\sqrt{n}}\right)$ for shallow networks. For deep ReLU networks, Chen et al. (2021) developed the bound of the order $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma^{2}}\sqrt{\frac{m}{n}}\right)$ via Rademacher complexity (Bartlett et al., 2017), suffering from exponential depth dependence, explicit width requirement and suboptimal $\sqrt{1/n}$ scaling. Taheri et al. (2025) improved the result to $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma^{2}n}\right)$ for deep networks. The dependence on $n,\gamma$ is optimal up to a logarithmic factor (Shamir, 2021; Schliserman and Koren, 2023). However, their results require smooth activations and exponential width in $L$. Our rate is almost-optimal and enjoys polynomial dependence over the network depth. Furthermore, our bound holds under the overparameterization $\widetilde{\Omega}(1/\gamma^{8})$, matching the requirement in Ji and Telgarsky (2020); Chen et al. (2021). This is much better than $1/\gamma^{6L+4}$ in Taheri et al. (2025). To the best of our knowledge, these are the best generalization bound and width condition for deep neural networks.
@@ 5 Conclusion and Future Work
There remain several interesting questions for future works. First, it is an interesting question to extend our methods to SGD. Second, while we establish polynomial Lipschitz constants near initialization, investigating whether similar bounds hold far from initialization would deepen our theoretical understanding. Finally, we only consider fully-connected neural networks. It is interesting to study the generalization analysis of networks with other architectures, such as CNNs and Resnets (Du et al., 2019b).
@@ Appendix A Technical Lemmas
We define the diagonal sign matrix $\Sigma^{l}(\mathbf{x})$ with $l\in[L]$ by
@@ A.1 Properties of the Initialization
Lemma 2 (Theorem 4.4.5 in Vershynin (2018)).
@@
Lemma 3 (Lemma 2.7.6 in Vershynin (2018)).
@@
Lemma 4 (Theorem 2.8.2 in Vershynin (2018)).
@@
$\|h_{0}^{l}(\mathbf{x}_{i})\|_{2}\in\left[\sqrt{\frac{2}{3}},\sqrt{\frac{4}{3}}\right],\quad i=1,\cdots,n\quad\text{and}\quad l=1,\cdots,L.$
@@
This result directly follows Corollary A.2 in Zou et al. (2018), and we give the proof here for completeness. Note that for $1\leq i\leq n,1\leq l\leq L$,
$\left\|h_{0}^{l}(\mathbf{x}_{i})\right\|_{2}^{2}$ $=\left\|\sqrt{\frac{2}{m}}\sigma\left(\mathbf{W}^{l}(0)h_{0}^{l-1}(\mathbf{x}_{i})\right)\right\|_{2}^{2}=\frac{1}{m}\sum_{r=1}^{m}2\sigma^{2}(\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle).$
Condition on $h_{0}^{l-1}(\mathbf{x}_{i})$, we have $\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle\sim\mathcal{N}(0,\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2})$, hence
$\mathbb{E}2\sigma^{2}(\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle)=2\left\|h_{0}^{l-1}(\mathbf{x}_{i})\right\|_{2}^{2}\mathbb{E}_{u\sim\mathcal{N}(0,1)}\sigma^{2}(u)=\left\|h_{0}^{l-1}(\mathbf{x}_{i})\right\|_{2}^{2}$ (18)
@@
$\|2\sigma^{2}(\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle)-\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2}\|_{\phi_{1}}$
$\leq$ $2\|2\sigma^{2}(\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle)\|_{\phi_{1}}\leq C\|\sigma(\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle)\|_{\phi_{2}}^{2}\leq C\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2},$
where the last inequality is due to (17). Let $X_{r}=2\sigma^{2}(\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle)-\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2},d_{r}=1/m,K=C\left\|h_{0}^{l-1}(\mathbf{x}_{i})\right\|_{2}^{2}$ and apply Lemma 4. We have for any $0\leq t\leq 1$,
$\mathbb{P}(|\|h_{0}^{l}(\mathbf{x}_{i})\|_{2}^{2}-\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2}|\leq Ct\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2}|h_{0}^{l-1}(\mathbf{x}_{i}))$
$=$ $\mathbb{P}\left(\left|\frac{1}{m}\sum_{r=1}^{m}(\mathbb{E}2\sigma^{2}(\langle w_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}_{i})\rangle)-\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2}\right|\leq Ct\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2}|h_{0}^{l-1}(\mathbf{x}_{i})\right)$
@@
$\mathbb{P}(|\|h_{0}^{l}(\mathbf{x}_{i})\|_{2}^{2}-\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2}|\leq Ct\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2})\geq 1-2nL\exp(-Cmt^{2}).$
@@
$|\|h_{0}^{l}(\mathbf{x}_{i})\|_{2}^{2}-\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2}|\leq C\sqrt{\frac{\log(nL/\delta)}{m}}\|h_{0}^{l-1}(\mathbf{x}_{i})\|_{2}^{2},\quad 1\leq i\leq n,1\leq l\leq L.$ (19)
@@
$|\|h_{0}^{l}(\mathbf{x}_{i})\|_{2}^{2}-1|\leq\frac{4lC}{3}\sqrt{\frac{\log(nL/\delta)}{m}}\leq\frac{1}{3}.$ (20)
When $l=0$, it is true. If (20) holds for $l\in[L-1]$, then $\|h_{0}^{l}(\mathbf{x}_{i})\|_{2}^{2}\leq 4/3$. Combined with (19), we have with probability at least $1-\delta$,
$|\|h_{0}^{l+1}(\mathbf{x}_{i})\|_{2}^{2}-1|$ $\leq|\|h_{0}^{l}(\mathbf{x}_{i})\|_{2}^{2}-1|+|\|h_{0}^{l+1}(\mathbf{x}_{i})\|_{2}^{2}-\|h_{0}^{l}(\mathbf{x}_{i})\|_{2}^{2}|$
$\leq\frac{4lC}{3}\sqrt{\frac{\log(nL/\delta)}{m}}+C\sqrt{\frac{\log(nL/\delta)}{m}}\|h_{0}^{l}(\mathbf{x}_{i})\|_{2}^{2}$
@@
$\left\|\frac{\partial f_{\mathbf{W}(0)}(\mathbf{x}_{i})}{\partial\mathbf{W}^{L}(0)}\right\|_{F}\leq\sqrt{2}.$
@@
By Hoeffding inequality, condition on $h_{0}^{L-1}(\mathbf{x}_{i})$, with probability at least $1-\delta$, there holds
$\frac{1}{m}\mathbf{a}^{\top}\Sigma^{L}_{0}(\mathbf{x}_{i})\Sigma_{0}^{L}(\mathbf{x}_{i})\mathbf{a}=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}\{\langle\mathbf{w}^{L}_{j}(0),h_{0}^{L-1}(\mathbf{x}_{i})\rangle\geq 0\}\leq\frac{1}{2}+C\sqrt{\frac{\log(n/\delta)}{m}}\leq\frac{3}{4}.$
@@
$\left\|\frac{\partial f_{\mathbf{W}(0)}(\mathbf{x}_{i})}{\partial\mathbf{W}^{L}(0)}\right\|_{F}^{2}=\frac{2}{m}\|h_{0}^{L-1}(\mathbf{x}_{i})\|_{2}^{2}\mathbf{a}^{\top}\Sigma^{L}_{0}(\mathbf{x}_{i})\Sigma_{0}^{L}(\mathbf{x}_{i})\mathbf{a}\leq 2\cdot\frac{4}{3}\cdot\frac{3}{4}=2.$
@@
$\underset{\|\Sigma_{1}\|_{0}\leq s}{\sup}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})\Sigma_{1}\|_{2}\lesssim 1.$ (21)
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})\Sigma_{1}v\|_{2}\lesssim 1.$ (22)
@@
$\underset{v\in\mathcal{P}}{\sup}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}_{i})v\|_{2}\lesssim 1.$ (23)
Let $\mathcal{S}$ be a subspace of $S^{m-1}$ that has at most $s$ non-zero coordinates. For such a subspace, we choose a $1/2$-cover of it and denote this cover by $\mathcal{Q}$. By Lemma 4.2.13 in Vershynin (2018),

### v2

@@ 4.2 Generalization Analysis
In previous work (chen2021overparameterizationsufficientlearndeep), they derived the bound of $\widetilde{O}\big(4^{L}L\sqrt{{mF(\overline{\mathbf{W}})}/{n}}\big)$. Specifically, they use the generalization analysis in bartlett2017spectrally, which requires to estimate the following term
@@
To bound the term $G^{\prime}=\sup_{z,\mathbf{W}\in\mathcal{W}_{1}}\ell(yf_{\mathbf{W}}(\mathbf{x}))$, we analyze the difference $f_{\mathbf{W}}(\mathbf{x})-f_{\overline{\mathbf{W}}}(\mathbf{x})$. Since both $\mathbf{W},\overline{\mathbf{W}}\in\mathcal{B}_{R}(\mathbf{W}(0))$ for some $R$, we only need to study the local variation $f_{\mathbf{W}}(\mathbf{x})-f_{\mathbf{W}(0)}(\mathbf{x})$. This approach necessitates characterizing the uniform behavior of deep networks in $\mathcal{B}_{R}(\mathbf{W}(0))$, specifically establishing control over their Lipschitz constants near initialization. Existing works usually lead to an exponential dependence on $L$ (JMLR:v25:23-0740; taheri2025sharper), thus resulting in a $e^{O(L)}$ term in the generalization bound. In particular, Lemma F.3 and F.5 in NEURIPS2020_b7ae8fec pointed out that $\|h^{l}(\mathbf{x})\|\leq C^{L},\|\partial f_{\mathbf{W}}(\mathbf{x})/\partial h^{l}(\mathbf{x})\|_{2}\leq C^{L-l+1}\sqrt{m}$. Based on these observations, taheri2025sharper showed that
@@
reducing our task to bounding the hidden layer perturbation. Previous approaches, including JMLR:v25:23-0740 and du2019gradientdescentprovablyoptimizes, employ a recursive estimation:
@@
In contrast to previous work, we develop the covering-number strategy to avoid the exponential dependence on depth. Specifically, we first show that for $n$ training data, if $m=\widetilde{\Omega}(L^{10}\log(n)R^{2})$, then $\|h^{l}(\mathbf{x}_{i})-h^{l}_{0}(\mathbf{x}_{i})\|_{2}=\widetilde{O}\left(\frac{L^{2}R}{\sqrt{m}}\right)$ holds for $i\in[n],l\in[L]$ (Lemma 15). We further take a $1/(C^{L}\sqrt{m})$-covering $D=\{\mathbf{x}^{j}:j=1,\ldots,|D|\}$ of the input space. Recall that the input space $\mathcal{X}=S^{d-1}$, it is well known from Corollary 4.2.13 in vershynin2018high that the number of $1/(C^{L}\sqrt{m})$-covering is given by $|D|\leq(1+2C^{L}\sqrt{m})^{d}$. Applying Lemma 15 to $D$ derives that if $m=\widetilde{\Omega}(L^{10}\log(|D|)R^{2})$
@@ 4.3 Optimal rates on NTK-separable data
In this section, we apply our general analysis to NTK-separable data (ji2020polylogarithmicwidthsufficesgradient; nitanda2020gradientdescentlearnoverparameterized; chen2021overparameterizationsufficientlearndeep; JMLR:v25:23-0422; deora2023optimization), and obtain the optimal rates.
@@
This means that the dataset is separable by the NTK feature at initialization with a margin $\gamma$. nitanda2020gradientdescentlearnoverparameterized pointed out that this assumption is weaker than positive eigenvalues of NTK Gram matrix, which has been widely used in the literature (du2019gradientdescentprovablyoptimizes; pmlr-v97-du19c; pmlr-v97-arora19a). With the above assumption, we have the following optimal risk bound on NTK separable data. The proof is given in Appendix D.
@@
Under Assumption 3, Theorem 3.3 and Proposition 4.2 of chen2021overparameterizationsufficientlearndeep show that when the network width satisfies $m=\widetilde{\Omega}(L^{22}/\gamma^{8})$, the training error is of the order $\widetilde{O}(1/T)$. We achieve the same guarantee under a significantly milder width condition of $m=\widetilde{\Omega}(L^{16}/\gamma^{8})$. This improvement is enabled by two key technical advances: a sharper bound for (4), and a tighter estimate of the iterate distance $\|\mathbf{W}(t)-\overline{\mathbf{W}}\|_{F}$. Specifically, we improve the bound in (4) by a factor of $L^{1/3}$, and show that $\|\mathbf{W}(t)-\overline{\mathbf{W}}\|_{F}=\widetilde{O}(1/\gamma)$, improving upon the previous $\widetilde{O}(\sqrt{L}/\gamma)$ bound. Together, these refinements reduce the required network width by a factor of $L^{6}$.
@@
Under similar assumptions, ji2020polylogarithmicwidthsufficesgradient derived the bound $\widetilde{O}\left(\frac{1}{\gamma^{2}\sqrt{n}}\right)$ for shallow networks, which was recently improved to $\widetilde{O}(\frac{1}{\gamma^{2}n})$ based on an improved control of the Rademacher complexity (lei2024optimization). For deep ReLU networks, chen2021overparameterizationsufficientlearndeep developed the bound of the order $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma}\sqrt{\frac{m}{n}}\right)$ via Rademacher complexity (bartlett2017spectrally), suffering from exponential depth dependence, explicit width requirement and suboptimal $\sqrt{1/n}$ scaling. taheri2025sharper improved the result to $\widetilde{O}\left(\frac{e^{O(L)}}{\gamma^{2}n}\right)$ for deep networks. The dependence on $n,\gamma$ is optimal up to a logarithmic factor (JMLR:v22:20-997; NEURIPS2023_d8ca28a3). However, their results require smooth activations and exponential width in $L$. Our rate is almost-optimal and enjoys polynomial dependence over the network depth. Furthermore, our bound holds under the overparameterization $\widetilde{\Omega}(1/\gamma^{8})$, matching the requirement in ji2020polylogarithmicwidthsufficesgradient; chen2021overparameterizationsufficientlearndeep. This is much better than $1/\gamma^{6L+4}$ in taheri2025sharper. To the best of our knowledge, these are the best generalization bound and width condition for deep neural networks.
5 Experiments
In this section, we make some experimental verifications to support our excess risk analysis. Our excess risk analysis in Theorem 3 imposes an NTK separability assumption, which has been validated in the literature. For example, (ji2020polylogarithmicwidthsufficesgradient) demonstrates that Assumption 3 holds for a noisy 2-XOR distribution, where the dataset is structured as follows:
$(x_{1},x_{2},y,\ldots,x_{d})\in$ $\left\{\left(\tfrac{1}{\sqrt{d-1}},0,1\right),\left(0,\tfrac{1}{\sqrt{d-1}},-1\right),\right$.
$\left.\left(-\tfrac{1}{\sqrt{d-1}},0,1\right),\left(0,-\tfrac{1}{\sqrt{d-1}},-1\right)\right\}\times\left\{-\tfrac{1}{\sqrt{d-1}},\tfrac{1}{\sqrt{d-1}}\right\}^{d-2}.$
Here, the factor $\frac{1}{\sqrt{d-1}}$ ensures that $\|x\|_{2}=1$, $\times$ above denotes the Cartesian product, and the label $y$ only depends on the first two coordinates of the input $x$. As shown in ji2020polylogarithmicwidthsufficesgradient, this dataset satisfies Assumption 3 with $1/\gamma=O(d)$, which implies that our excess risk bound in Theorem 3 becomes $O(d^{2}/n)$ for this dataset. We conducted numerical experiments and observed that the test error decays linearly with $d^{2}/n$. The population loss for the test error is computed over all $2^{d}$ points in the distribution.
Settings
We train two-layer ReLU networks by gradient descent on noisy 2-XOR data. We fix the width $m=128,T=500,\eta=0.1$. We have conducted two experiments. With a fixed dimension $d=6$, we vary the sample size $n$. The results are presented in Table 2. With a fixed sample size $n=64$, we vary the dimension $d$ and the corresponding table is provided in Table 3.
Table 2: Test error for different $n$
$n$ 10 12 14 16 18 20 24 28
$d^{2}/n$ 3.60 3.00 2.57 2.25 2.00 1.80 1.50 1.29
Test Error 0.4625 0.4500 0.3625 0.3438 0.3219 0.3063 0.2469 0.2062
Table 3: Test error for different $d$
$d$ 7 8 9 10 11 12
$d^{2}/n$ 0.77 1.00 1.27 1.56 1.89 2.25
Test Error 0.0125 0.1313 0.2484 0.3080 0.3365 0.4190
In both experiments, we observe that the test error is of the order $d^{2}/n$ (approximately $0.15d^{2}/n$). This shows the consistency between our excess risk bounds in Theorem 3 and experimental results. We conducted the experiments on Google Colab. A simple demonstration reproducing our numerical experiments is available as a Google Colab notebook at: https://github.com/YuanfanLi2233/nips2025-optimal.
@@ 6 Conclusion and Future Work
There remain several interesting questions for future works. First, it is an interesting question to extend our methods to SGD. Second, while we establish polynomial Lipschitz constants near initialization, investigating whether similar bounds hold far from initialization would deepen our theoretical understanding. Finally, we only consider fully-connected neural networks. It is interesting to study the generalization analysis of networks with other architectures, such as CNNs and Resnets (du2019gradientdescentprovablyoptimizes).
Acknowledgement
The authors are grateful to the anonymous reviewers for their thoughtful comments and constructive suggestions. The work of Yuanfan Li and Zheng-Chu Guo is partially supported by the National Natural Science Foundation of China (Grants No. 12271473 and U21A20426). The work by Yunwen Lei is partially supported by the Research Grants Council of Hong Kong [Project No. 17302624]. Yiming’s work is partially supported by Australian Research Council (ARC) DP250101359.
@@ Appendix A Technical Lemmas
Given a set of $N$ points on the sphere $K=\{x^{1},\cdots,x^{N}\}$. We provide general results for any finite set $K$, then it can be applied to specific choices of $K$, for example, the training dataset $S_{1}=\{\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\}$. We define the diagonal sign matrix $\Sigma^{l}(\mathbf{x})$ with $l\in[L]$ by
@@ A.1 Properties of the Initialization
Lemma 2 (Theorem 4.4.5 in vershynin2018high).
@@
Lemma 3 (Lemma 2.7.6 in vershynin2018high).
@@
Lemma 4 (Theorem 2.8.2 in vershynin2018high).
@@
$\|h_{0}^{l}(\mathbf{x}^{i})\|_{2}\in\left[\sqrt{\frac{2}{3}},\sqrt{\frac{4}{3}}\right],\quad i=1,\cdots,N\quad\text{and}\quad l=1,\cdots,L.$
@@
This result directly follows Corollary A.2 in zou2018stochasticgradientdescentoptimizes, and we give the proof here for completeness. Note that for $1\leq i\leq N,1\leq l\leq L$,
$\left\|h_{0}^{l}(\mathbf{x}^{i})\right\|_{2}^{2}$ $=\left\|\sqrt{\frac{2}{m}}\sigma\left(\mathbf{W}^{l}(0)h_{0}^{l-1}(\mathbf{x}^{i})\right)\right\|_{2}^{2}=\frac{1}{m}\sum_{r=1}^{m}2\sigma^{2}(\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle).$
Condition on $h_{0}^{l-1}(\mathbf{x}^{i})$, we have $\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle\sim\mathcal{N}(0,\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2})$, hence
$\mathbb{E}2\sigma^{2}(\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle)=2\left\|h_{0}^{l-1}(\mathbf{x}^{i})\right\|_{2}^{2}\mathbb{E}_{u\sim\mathcal{N}(0,1)}\sigma^{2}(u)=\left\|h_{0}^{l-1}(\mathbf{x}^{i})\right\|_{2}^{2}$ (18)
@@
$\|2\sigma^{2}(\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle)-\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2}\|_{\phi_{1}}$
$\leq$ $2\|2\sigma^{2}(\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle)\|_{\phi_{1}}\leq C\|\sigma(\langle\mathbf{w}_{j}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle)\|_{\phi_{2}}^{2}\leq C\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2},$
where the last inequality is due to (17). Let $X_{r}=2\sigma^{2}(\langle\mathbf{w}_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle)-\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2},d_{r}=1/m,K=C\left\|h_{0}^{l-1}(\mathbf{x}^{i})\right\|_{2}^{2}$ and apply Lemma 4. We have for any $0\leq t\leq 1$,
$\mathbb{P}(|\|h_{0}^{l}(\mathbf{x}^{i})\|_{2}^{2}-\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2}|\leq Ct\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2}|h_{0}^{l-1}(\mathbf{x}^{i}))$
$=$ $\mathbb{P}\left(\left|\frac{1}{m}\sum_{r=1}^{m}(\mathbb{E}2\sigma^{2}(\langle w_{r}^{l}(0),h_{0}^{l-1}(\mathbf{x}^{i})\rangle)-\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2}\right|\leq Ct\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2}|h_{0}^{l-1}(\mathbf{x}^{i})\right)$
@@
$\mathbb{P}(|\|h_{0}^{l}(\mathbf{x}^{i})\|_{2}^{2}-\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2}|\leq Ct\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2})\geq 1-2nL\exp(-Cmt^{2}).$
@@
$|\|h_{0}^{l}(\mathbf{x}^{i})\|_{2}^{2}-\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2}|\leq C\sqrt{\frac{\log(NL/\delta)}{m}}\|h_{0}^{l-1}(\mathbf{x}^{i})\|_{2}^{2},\quad 1\leq i\leq N,1\leq l\leq L.$ (19)
@@
$|\|h_{0}^{l}(\mathbf{x}^{i})\|_{2}^{2}-1|\leq\frac{4lC}{3}\sqrt{\frac{\log(NL/\delta)}{m}}\leq\frac{1}{3}.$ (20)
When $l=0$, it is true. If (20) holds for $l\in[L-1]$, then $\|h_{0}^{l}(\mathbf{x}^{i})\|_{2}^{2}\leq 4/3$. Combined with (19), we have with probability at least $1-\delta$,
$|\|h_{0}^{l+1}(\mathbf{x}^{i})\|_{2}^{2}-1|$ $\leq|\|h_{0}^{l}(\mathbf{x}^{i})\|_{2}^{2}-1|+|\|h_{0}^{l+1}(\mathbf{x}^{i})\|_{2}^{2}-\|h_{0}^{l}(\mathbf{x}^{i})\|_{2}^{2}|$
$\leq\frac{4lC}{3}\sqrt{\frac{\log(NL/\delta)}{m}}+C\sqrt{\frac{\log(NL/\delta)}{m}}\|h_{0}^{l}(\mathbf{x}^{i})\|_{2}^{2}$
@@
$\left\|\frac{\partial f_{\mathbf{W}(0)}(\mathbf{x}^{i})}{\partial\mathbf{W}^{L}(0)}\right\|_{F}\leq\sqrt{2}.$
@@
By Hoeffding inequality, condition on $h_{0}^{L-1}(\mathbf{x}^{i})$, with probability at least $1-\delta$, there holds
$\frac{1}{m}\mathbf{a}^{\top}\Sigma^{L}_{0}(\mathbf{x}^{i})\Sigma_{0}^{L}(\mathbf{x}^{i})\mathbf{a}=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}\{\langle\mathbf{w}^{L}_{j}(0),h_{0}^{L-1}(\mathbf{x}^{i})\rangle\geq 0\}\leq\frac{1}{2}+C\sqrt{\frac{\log(N/\delta)}{m}}\leq\frac{3}{4}.$
@@
$\left\|\frac{\partial f_{\mathbf{W}(0)}(\mathbf{x}^{i})}{\partial\mathbf{W}^{L}(0)}\right\|_{F}^{2}=\frac{2}{m}\|h_{0}^{L-1}(\mathbf{x}^{i})\|_{2}^{2}\mathbf{a}^{\top}\Sigma^{L}_{0}(\mathbf{x}^{i})\Sigma_{0}^{L}(\mathbf{x}^{i})\mathbf{a}\leq 2\cdot\frac{4}{3}\cdot\frac{3}{4}=2.$
@@
$\underset{\|\Sigma_{1}\|_{0}\leq s}{\sup}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})\Sigma_{1}\|_{2}\lesssim 1.$ (21)
@@
$\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})\Sigma_{1}v\|_{2}\lesssim 1.$ (22)
@@
$\underset{v\in\mathcal{P}}{\sup}\|\mathbf{H}_{b,0}^{a}(\mathbf{x}^{i})v\|_{2}\lesssim 1.$ (23)
Let $\mathcal{S}$ be a subspace of $S^{m-1}$ that has at most $s$ non-zero coordinates. For such a subspace, we choose a $1/2$-cover of it and denote this cover by $\mathcal{Q}$. By Lemma 4.2.13 in vershynin2018high,